OpenStudy (anonymous):

Determine the length of the hypotenuse in a right triangle with legs 5 inches and 12 inches.

OpenStudy (anonymous):

PLease help @L-Lawliet-L THNKS SOOO MCH!

OpenStudy (anonymous):

a^2=b^2+b^2 a^2=hypotenuse

OpenStudy (anonymous):

a^1=25+144 a^2=169 a=13

OpenStudy (ivettef365):

use Pythagorean theorem as L-Lasliet formula shown